A card is drawn from a deck, and a game spinner with three equal regions labeled 1, 2, and 3 is spun. Let A be the event that the number on the card is 2. Let B be the event that the number on the spinner is a 2.

a.Find the number of elements in the sample space.
b.Find the number of elements in the set describing each event: A, B, A union B, A intersection B. And find the probability of each.

Answer by stanbon(75887)

A card is drawn from a deck, and a game spinner with three equal regions labeled 1, 2, and 3 is spun.
Let A be the event that the number on the card is 2.
Let B be the event that the number on the spinner is a 2.
a.Find the number of elements in the sample space: 52*3 = 156
-------------------
b.Find the number of elements in the set describing each event:
And find the probability of each.
---------------
A: 4*3 = 12 ; P(A) = 12/156
B: (1/3)156 = 52
A union B: 12+52-4 = 60 ; P(AUB) = 60/156
A intersection B: 4 : P(A and B) = 4/156
==============================================
Cheers,
Stan H.
============